OpenStudy (anonymous):

how do you reduce 22 2/4

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

22 and 2/4 = \(22 + \frac{2}{4} = \frac{88}{4} + \frac{2}{4} = \frac{90}{4} = \frac{45*2}{2*2} = \frac{45}{2}\)

OpenStudy (anonymous):

so would it be 45 and 1/7 or 45 1/6

OpenStudy (anonymous):

Uh.. no. \(\frac{45}{2}\) = 22 and 1/2. I didn't realize you wanted it in compound form. Otherwise I woulda just reduced the 2/4 to 1/2 and called it a day.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

1 more thing what is the answer to this it took smedley 5 3/4 hours to climb the top of the mountain it took 3 1/4 hours to climb down if he spent 1 1/2 hours at the top how long did the climb take? And how